# STANCE-$\text{C}^{3}$: Domain-adaptive Cross-target Stance Detection via Contrastive Learning and Counterfactual Generation

### v1

@@ III-B Domain-adaptive Cross-target
One of the main challenges in detecting stances across different targets is that the distribution of data varies for each target even within the same domain. For instance, the word WHO may appear more frequently in the domain where the target is COVID-19 vaccination than in another domain where the target is Donald Trump [13]. The common occurrences between specific target words and instances make the model biased toward learning. Therefore it is necessary to identify an effective way to model transferable knowledge. To learn target-invariant features in instances from both the origin $t^{o}$ and destination $t^{d}$ targets, we use a simple yet effective supervised contrastive learning approach.
Contrastive learning has been widely used for many applications in Computer Vision, such as identity recognition tasks. The goal of contrastive learning is to pull the distance between an anchor and positive pair and push the distance between an anchor and negative pair. Here, we use the supervised contrastive loss [14], which applies a modified version of contrastive learning to supervised learning settings. The supervised contrastive loss function aims to enhance the separability of different classes by maximizing the distance between the anchor and the negative samples by pushing away the samples from different classes while minimizing the distance between the anchor and the positive samples by pulling together the data points from the same class.
@@
In our problem setting, considering samples with different stance targets $t^{o}$ and $t^{d}$, the goal of the contrastive learning approach is to reduce the distance between samples with the same stance label and vice-versa. For example, we aim to maximize the distance between pairs of representations obtained from tweets that exhibit different stances towards the same target. Note that contrastive learning is applied to all pairs, regardless of whether the targets in the pair are equivalent or not. This approach creates a representation for the stance classifier that indicates the relation between the statement and the stance target $t$. We use the following modified supervised contrastive learning method in the spirit of [14]:
@@
The loss takes a batch of samples $i\in I$ as an "anchor" from the training set where each sample is mapped to a vector $z_{i}$. Each anchor allows for multiple positive vectors $z_{p}$ that have the same label as an anchor ($y_{j}=y_{i}$) in addition to multiple negative vectors $z_{n}$ ($z_{n}\in A(i)$, $y_{n}\neq y_{i}$). $P(i)$ and $N(i)$ represent the indices of all positive and negative vectors, respectively. $\tau$ shows the temperature parameter that affects the distance of instance discrimination. The notation $sim(\cdot)$ refers to the cosine similarity. Here, we introduce the second term to balance the weight of positive and negative pairs. According to our observations on experiments involving different variants of the modified contrastive loss, we discovered that weighing the original loss with text similarity improves the model’s generalization ability across different targets. Note that anchor $z_{i}$ is from target $t^{o}$ but $z_{p}$ and $z_{n}$ are from both $t^{d}$ and $t^{o}$ for target-agnostic feature learning.
@@
Optimization Algorithm The training process of STANCE-$\text{C}^{3}$ has two stages [^1] Source code and datasets will become publicly available upon acceptance.. During the first stage, the parameters of the T5-based LM are trained for conversion between source $\mathcal{S}$ and the target $\mathcal{T}$ domain following the work of Calderon et al. [12]. This stage enriches the input dataset to include more samples similar to the target domain. The next stage uses the contrastive learning approach to create a cross-target sentence representation. During the contrastive learning process, we add an additional loss $\ell_{\mathrm{cont}}$ to the classification loss for cross-target stance detection. For each instance $x_{i}$, the overall loss is formulated as $\ell_{\mathrm{total}}=\lambda\ell_{\mathrm{cont}}+(1-\lambda)\ell_{\mathrm{CE}}$ where $\ell_{\mathrm{CE}}$ stands for the cross-entropy loss for classification and $\lambda$ is the weighting factor. The cross-entropy loss is calculated as $CE=-\sum_{i=1}^{k}y_{i}\log(p_{i})$ for $k$ classes based on the true stance label $y$ and predicted stance probability $p$ for the $i^{th}$ class. Similar to [14] we set the temperature parameter $\tau=0.08$.
IV Experiments
@@
The experiments conducted in this study investigate how changing the domains and targets affect the performance of several stance detection models, including STANCE-$\text{C}^{3}$. We also examine the effects of data augmentation and contrastive learning on performance accuracy in cross-domain and cross-target scenarios. More specifically, we aim to answer the following research questions: Q1. Does STANCE-$\text{C}^{3}$  achieve better performance in comparison to other SOTA when trained on source domain $\mathcal{S}$ with some data from target domain $\mathcal{T}$, and tested on target domain $\mathcal{T}$ for target words in the source domain $\mathcal{S}$? Q2. Does STANCE-$\text{C}^{3}$  achieve better performance in comparison to other SOTA when trained on a source domain $\mathcal{S}$ and tested on a target domain $\mathcal{T}$ for target words not in the source domain? Q3. What is the effect of changes in the model’s parameters and components on its performance in the stance detection task? We consider the following scenarios in the experimental design: (1) single domain + single target, (2) cross-domain + single target, and (3) cross-domain + cross-target, where the targets and domains of the datasets have clear distributional differences. We address Q1 by training and testing the baseline and proposed models under the first scenario where both domain and target are the same for training and testing sets. For Q2, the models are evaluated under scenarios (2) and (3) to compare and contrast the models’ domain-adaptive and cross-target performance. Finally, we address Q3 by performing ablation study and parameter analysis on datasets with different stance targets.
@@
TABLE II: Baselines targeted goals indicate the difference between the proposed approach STANCE-$\text{C}^{3}$ and the baselines.
@@
IV-A Baselines
We consider the following baselines. Table II shows the difference between baselines and the proposed approach. Note that we have implemented the baselines for both cross-domain and cross-target tasks in the table, regardless of their specific task targets. This choice stems from the unique nature of the task, which presents limitations in terms of available baselines.
RLFND [16]: A domain-adaptive reinforcement learning framework for fake news detection. The proposed method leverages a deep reinforcement learning algorithm to learn the optimal policy for fake news detection while adapting to the domain shift between the source and target domains.
MTSD [17]: A multi-task learning framework that performs stance detection on multiple targets simultaneously. The framework leverages a shared encoder to capture the common features of the text and a task-specific decoder to predict the stance toward each target.
We compare our approach to the existing state-of-the-art approaches using accuracy and AUC metrics.
IV-B Datasets
TABLE III: Datasets’ statistics.
Dataset Target Source Labels Samples
CoVaxNet
COVID-19 Vaccination
Twitter
Pro, Anti
1,831,220
COVID-19-Stance
Face masks, Fauci, School closures, Stay at home orders
Twitter
Favor, Against, None
7,122
We use two representative datasets related to COVID-19 with different domains and targets (see Table III). The large-scale Twitter dataset CoVaxNet [18] was used for both cross-domain and cross-target training. CoVaxNet was acquired over the pandemic period (Jan 1st, 2020 - Dec 31st, 2021) and labeled with stances toward “COVID-19 vaccination”. For testing the performance of STANCE-$\text{C}^{3}$  in a single-target setting, the domains are defined as separate time periods marked by a changepoint event. We divided CoVaxNet into two time periods with the changepoint event occurring on August 23, 2021, when the US Food and Drug Administration (FDA) fully approved the Pfizer-BioNTech vaccine. Thus, the data for the source and target domains are 10,000 randomly selected tweets from Jan 1st, 2020 to August 22, 2021 (“CoVaxNet_pre”; source domain) and 10,000 tweets from August 23, 2021 to December 31st, 2021 (“CoVaxNet_post”; target domain). For the cross-target scenario, CoVaxNet_pre and CoVaxNet_post were separately used as training sets for STANCE-$\text{C}^{3}$. Model performance was tested on targets (“Anthony S. Fauci, M.D.”, “Wearing a Face Mask”, “Keeping Schools Closed”, “Stay at Home Orders”) from the COVID-19-Stance [19] dataset. COVID-19-Stance was collected in the period (February 27th, 2020 - August 20th, 2020) with different targets and keywords of interest (e.g., #lockdown), thus justifying an adequate distributional shift between the source and target. It should be noted that we treated the congruent labels from the two datasets as equivalent i.e., pro is equivalent to favor and anti is the same as against. In the context of stance detection, the words in these pairs essentially convey the same stance.
TABLE IV: Cross-domain, single-target performance results. In this experiment we only use the datasets that have the same target word (p-value < 0.05 for all McNemar’s tests).
Source Domain Target Domain BERT MoLE MTL MTSD RLFND STANCE-$\text{C}^{3}$
CoVaxNet_pre CoVaxNet_post 0.713 (0.697) 0.646 (0.621) 0.734 (0.736) 0.734 (0.731) 0.752 (0.747) 0.765 (0.753)
CoVaxNet_pre 0.762 (0.727) 0.901 (0.876) 0.823 (0.789) 0.845 (0.831) 0.843 (0.835) 0.863 (0.858)
Performance Degradation 0.049 (0.030) 0.255 (0.255) 0.089 (0.053) 0.111 (0.100) 0.091 (0.088) 0.098 (0.105)
CoVaxNet_post CoVaxNet_post 0.826 (0.796) 0.821 (0.835) 0.811 (0.823) 0.841 (0.839) 0.812 (0.784) 0.831 (0.837)
CoVaxNet_pre 0.698 (0.655) 0.713 (0.702) 0.623 (0.611) 0.712 (0.692) 0.674 (0.632) 0.723 (0.723)
Performance Degradation 0.128 (0.141) 0.108 (0.133) 0.188 (0.212) 0.129 (0.147) 0.138 (0.152) 0.108 (0.114)
TABLE V: Cross-domain, cross-target performance results. In this experiment, we evaluate the model’s performance on a dataset that has a different target word in comparison to the source domain’s target (p-value < 0.05 for all McNemar’s tests).
Source Domain Target Domain BERT MoLE MTL MTSD RLFND STANCE-$\text{C}^{3}$
CoVaxNet_pre Face Masks 0.623 (0.568) 0.732 (0.711) 0.741 (0.712) 0.789 (0.746) 0.742 (0.625) 0.868 (0.823)
Fauci 0.741 (0.706) 0.725 (0.719) 0.691 (0.687) 0.732 (0.698) 0.699 (0.625) 0.743 (0.714)
School Closures 0.832 (0.627) 0.778 (0.761) 0.711 (0.651) 0.872 (0.731) 0.834 (0.804) 0.920 (0.734)
Stay at Home 0.641 (0.625) 0.734 (0.738) 0.621 (0.613) 0.738 (0.732) 0.665 (0.658) 0.736 (0.742)
Average Performance 0.709 (0.631) 0.742 (0.732) 0.691 (0.665) 0.782 (0.726) 0.735 (0.678) 0.816 (0.753)
CoVaxNet_post Face Masks 0.570 (0.555) 0.801 (0.786) 0.698 (0.676) 0.794 (0.754) 0.734 (0.712) 0.819 (0.769)
Fauci 0.548 (0.624) 0.721 (0.740) 0.709 (0.683) 0.763 (0.778) 0.713 (0.694) 0.743 (0.714)
School Closures 0.578 (0.543) 0.718 (0.743) 0.621 (0.620) 0.687 (0.701) 0.694 (0.675) 0.739 (0.772)
Stay at Home 0.623 (0.637) 0.672 (0.691) 0.669 (0.681) 0.699 (0.704) 0.675 (0.662) 0.725 (0.727)
Average Performance 0.579 (0.589) 0.728 (0.740) 0.674 (0.665) 0.735 (0.734) 0.704 (0.685) 0.756 (0.745)
IV-C Evaluation and Results
Table IV and Table V summarize the comparative performance of the baseline and proposed models with respect to the accuracy and AUC (in parentheses).
Q1: In this experiment, the proposed approach and the domain-adaptive baselines used a fixed portion (30%) of the target domain data during counterfactual data generation. The models were trained on a single domain and tested on both source and target domains for stance detection for a single target of interest from the source domain. Our results provide some evidence that even when the target of interest is contained in the source domain, performance in the target domain generally degrades for all models. The Performance Degradation in the table highlights how a model’s performance changes when shifting from the source domain to the target domain. We observe that all of the models exhibit performance degradation which confirms the difference between domains.
Q2: The results of our experiments on the cross-domain cross-target setting, as presented in Table V, demonstrate the performance of our approach and the baselines. In this setting, we augment the training set of baselines with a small portion (30%) of data from the target domain. The results of our experiments suggest that the proposed approach is more effective than the domain-adaptive baselines. Specifically, our approach surpasses all baselines in terms of average accuracy and AUC score, underscoring its suitability for the cross-domain cross-target context. Furthermore, these results underscore the robustness of our approach to variations in the target domain, a strength attributable to our counterfactual data generation technique which is tailored to produce data that better represents the target domain.
Parameter Analysis To answer Q3, we analyze the effect of hyperparameters on the performance of STANCE-$\text{C}^{3}$. We show the effect of crucial training parameters $\lambda$ and the target domain portion $\gamma$ in the training set. To demonstrate the effect of contrastive loss and the balance between contrastive and cross-entropy loss, we vary the $\lambda$ value in the range of $\{0.0,0.25,0.5,0.75,1.0\}$. As illustrated in Figure 2-a, the AUC score varies for different $\lambda$ values. The results indicate that using $\lambda=0.5$ yields the best performance.
Next, we investigate the effect of the target domain data portion $\gamma$ used for counterfactual data generation. We vary the $\gamma$ value within the range of $\{5\%,15\%,30\%,45\%\}$. The results, as depicted in Figure 2-b, show that adding more target domain data leads to better performance. However, in real-world scenarios, access to a large amount of data in the target domain is often limited. Furthermore, there is no significant performance difference between $\gamma=30\%$ and $\gamma=45\%$. Therefore, we conclude that using $30\%$ of the target domain data enables STANCE-$\text{C}^{3}$ to achieve an acceptable performance.
[…]
V Related Work
STANCE-$\text{C}^{3}$ is proposed in this paper to improve existing models for stance detection and robustness across domains. In this section, we review current NLP SOTA models for stance detection and domain adaptation.
V-A Stance Detection
There are two general categories of distinctions among stance detection models. Firstly, models are distinguished on how stance features are modeled. Content-level approaches use linguistic features such as topics [20] or targets [21] alongside sentiment information [22]. User-level approaches focus on user-related attributes interactions [23], information [24], and timelines [25]. Hybrid models combine content- and user-level features for more comprehensive representations [26]. In this paper, we focus on content-level modeling with target embedding using BERT due to its practicality in data acquisition.
Secondly, stance detection models differ with respect to the specificity of the target of interest. While many previous works focus on specific single-target scenarios [27], recent studies have considered cross-target [21], multi-target [17], and few-shot or zero-shot stance detection [28]. While both cross-target and zero-shot stance detection models are trained on one or more source targets and used on previously unseen targets [29], zero-shot methods are able to predict stances for unseen targets that are not necessarily related to the target in the training set [30]. Both attempt to extract transferable knowledge from the source using methods such as hierarchical contrastive learning[31]. STANCE-$\text{C}^{3}$uses a similar but modified approach to the contrastive loss during training.
V-B Domain Adaptation
As a category of transfer learning, domain Adaptation (DA) leverages on agnostic or common information between domains with the target task remaining constant [32]. There are several DA setups, including unsupervised domain adaptation (UDA) where both labeled source data and unlabeled target data are available [33], semi-supervised domain adaptation (SSDA) where labeled source data and a small number of labeled target data are accessible [34], and supervised domain adaptation (SDA) where both labeled source and target data are available during training [35]. Our focus is on SSDA, where a smaller set of labeled target data is used in the training process.

### v2

@@ IV-B Domain-adaptive Cross-target
One of the main challenges in detecting stances across different targets is the variation in data distribution for each target, even within the same domain. For instance, the word WHO may appear more frequently in the domain where the target is COVID-19 vaccination than in another domain where the target is Donald Trump [39]. The frequent co-occurrences of specific target words with certain instances make the model biases the model’s learning process. Therefore it is necessary to identify an effective way to model transferable knowledge. To learn target-invariant features in instances from both the origin $t^{o}$ and destination $t^{d}$ targets, we use a simple yet effective supervised contrastive learning  [40] approach. This loss function aims to enhance the separability of different classes by maximizing the distance between an anchor and negative samples.
@@
In our problem setting, considering samples with different stance targets $t^{o}$ and $t^{d}$, the goal is to reduce the distance between samples sharing the same stance label and increase it for those with differing labels. Note that contrastive learning is applied to all pairs, regardless of whether the targets in the pair are equivalent or not. This approach creates a representation for the stance classifier that indicates the relation between the statement and the stance target $t$. We introduce a modified version of the supervised contrastive learning tailored to better address the nuances of our problem setting.
@@
The loss takes a batch of samples $i\in I$ as an “anchor” from the training set where each sample is mapped to a vector $z_{i}$. Each anchor allows for multiple positive vectors $z_{p}$ that have the same label as an anchor ($y_{j}=y_{i}$) in addition to multiple negative vectors $z_{n}$ ($z_{n}\in A(i)$, $y_{n}\neq y_{i}$). $P(i)$ and $N(i)$ represent the indices of all positive and negative vectors, respectively. $\tau$ shows the temperature parameter that affects the distance of instance discrimination. The notation $sim(\cdot)$ refers to the cosine similarity. Here, we introduce the second term to balance the weight of positive and negative pairs. According to our observations on experiments involving different variants of the modified contrastive loss, we discovered that weighing the original loss with text similarity improves the model’s generalization ability across different targets. Note that anchor $z_{i}$ is from target $t^{o}$ but $z_{p}$ and $z_{n}$ are from both $t^{d}$ and $t^{o}$ for target-agnostic feature learning.
@@
Optimization Algorithm The training process of STANCE-C3 for domain-adaptive cross-target scenario has two stages [^1] Source code and data will become publicly available upon acceptance. Partial source code and data are available at https://osf.io/26rfc/?view_only=b6e3bfac71444eb0be0aaad0f671a720. In the first stage, the parameters of the T5-based LM are trained for conversion between source $\mathcal{S}$ and the target $\mathcal{T}$ domain following the work of [38]. This stage enriches the input dataset to include more samples similar to the target domain. The next stage uses the contrastive learning approach to create a cross-target sentence representation. We randomly selected 50 anchors from $\mathcal{D}_{\mathcal{S}}$ and its domain counterfactuals, along with their positive and negative pairs from across all dataset except for generated counterfactuals. For each instance $x_{i}$, the overall loss is formulated as $\ell_{\mathrm{total}}=\lambda\ell_{\mathrm{cont}}+(1-\lambda)\ell_{\mathrm{CE}}$ where $\ell_{\mathrm{CE}}$ stands for the cross-entropy loss for classification and ($\ell_{\mathrm{cont}}$) represents the contrastive loss. $\lambda$ determines the balance between the two loss components. Note that while the classification loss is computed over the entire training dataset, the contrastive loss is derived solely from the selected pairs mentioned previously. The cross-entropy loss is calculated as $CE=-\sum_{i=1}^{k}y_{i}\log(p_{i})$ for $k$ classes based on the true stance label $y$ and predicted stance probability $p$ for the $i^{th}$ class. Similar to [40] we set the temperature parameter $\tau=0.08$.
V Experiments
TABLE I: Datasets’ statistics: For COVID-19-Stance dataset, a significant number of tweets could not be downloaded due to Twitter restrictions and deletion of the tweets. In this context, the number of samples mentioned on the left refers to the actual sample size whereas the original data count is on the right.
Dataset Target Platform Labels
CoVaxNet COVID-19 Vaccination Twitter Pro (1,495,991), Anti (335,229)
COVID-19-Stance Wearing a face mask, Anthony S. Fauci, M.D., Keeping schools closed, Stay at home orders Twitter In-favor (2,421/43,799), Against (1,577/29,693)
@@ V Experiments
The experiments conducted in this study investigate how changing the domains and targets affects the performance of several stance detection models, including STANCE-C3. We also examine the effects of data augmentation and contrastive learning on performance accuracy in cross-domain and cross-target scenarios. More specifically, we aim to answer the following research questions: Q1. Does STANCE-C3  achieve better performance in comparison to other SOTA when trained on source domain $\mathcal{S}$ with some data from target domain $\mathcal{T}$, and tested on target domain $\mathcal{T}$ for target words in the source domain $\mathcal{S}$? Q2. Does STANCE-C3  achieve better performance in comparison to other SOTA when trained on a source domain $\mathcal{S}$ and tested on a target domain $\mathcal{T}$ for target words not in the source domain? Q3. What is the effect of changes in the model’s parameters and components on its performance in the stance detection task? We consider the following scenarios in the experimental design: (1) single domain + single target, (2) cross-domain + single target, and (3) cross-domain + cross-target, where the targets and domains of the datasets have clear distributional differences. We address Q1 by training and testing the baseline and proposed models under scenarios (1) and (2) where targets are the same for training and testing sets. For Q2, the models are evaluated under scenario (3) to compare and contrast the models’ domain-adaptive and cross-target performance. Finally, we address Q3 by performing ablation studies and parameter analysis. Notably, domain counterfactual generation is applied in all scenarios, and contrastive learning is employed only for scenario (3).
@@
TABLE II: Baselines targeted goals indicate the difference between the proposed approach STANCE-C3 and the baselines.
@@
Llama-2-7B [42]
STANCE-C3
TABLE III: Cross-domain, single-target performance results in accuracy and AUC (in parentheses). In this experiment we only use the datasets that have the same target word (p-value $<$ 0.05 for all McNemar’s tests).
Source Domain Target Domain BERT MoLE MTL MTSD RLFND Llama-2-7B STANCE-C3
CoVaxNet_pre CoVaxNet_post 0.713 (0.697) 0.646 (0.621) 0.734 (0.736) 0.734 (0.731) 0.752 (0.747) 0.721 (0.725) 0.765 (0.753)
CoVaxNet_pre 0.762 (0.727) 0.901 (0.876) 0.823 (0.789) 0.845 (0.831) 0.843 (0.835) 0.812 (0.807) 0.863 (0.858)
Performance Degradation 0.049 (0.030) 0.255 (0.255) 0.089 (0.053) 0.111 (0.100) 0.091 (0.088) 0.091 (0.082) 0.098 (0.105)
CoVaxNet_post CoVaxNet_post 0.826 (0.796) 0.821 (0.835) 0.811 (0.823) 0.841 (0.839) 0.812 (0.784) 0.761 (0.719) 0.831 (0.837)
CoVaxNet_pre 0.698 (0.655) 0.713 (0.702) 0.623 (0.611) 0.712 (0.692) 0.674 (0.632) 0.688 (0.641) 0.723 (0.723)
Performance Degradation 0.128 (0.141) 0.108 (0.133) 0.188 (0.212) 0.129 (0.147) 0.138 (0.152) 0.073 (0.078) 0.108 (0.114)
TABLE IV: Cross-domain, cross-target performance results in accuracy and AUC (in parentheses). In this experiment, we evaluate the model’s performance on a dataset that has a different target word in comparison to the source domain’s target (p-value $<$ 0.05 for all McNemar’s tests).
Source Domain Target Domain BERT MoLE MTL MTSD RLFND Llama-2-7B STANCE-C3
CoVaxNet_pre Face Masks 0.623 (0.568) 0.732 (0.711) 0.741 (0.712) 0.789 (0.746) 0.742 (0.625) 0.862 (0.815) 0.868 (0.823)
Fauci 0.741 (0.706) 0.725 (0.719) 0.691 (0.687) 0.732 (0.698) 0.699 (0.625) 0.673 (0.635) 0.743 (0.714)
School Closures 0.832 (0.627) 0.778 (0.761) 0.711 (0.651) 0.872 (0.731) 0.834 (0.804) 0.894 (0.796) 0.920 (0.734)
Stay at Home 0.641 (0.625) 0.734 (0.738) 0.621 (0.613) 0.738 (0.732) 0.665 (0.658) 0.761 (0.759) 0.736 (0.742)
Average Performance 0.709 (0.631) 0.742 (0.732) 0.691 (0.665) 0.782 (0.726) 0.735 (0.678) 0.797 (0.751) 0.816 (0.753)
CoVaxNet_post Face Masks 0.570 (0.555) 0.801 (0.786) 0.698 (0.676) 0.794 (0.754) 0.734 (0.712) 0.808 (0.758) 0.819 (0.769)
Fauci 0.548 (0.624) 0.721 (0.740) 0.709 (0.683) 0.763 (0.778) 0.713 (0.694) 0.678 (0.644) 0.743 (0.714)
School Closures 0.578 (0.543) 0.718 (0.743) 0.621 (0.620) 0.687 (0.701) 0.694 (0.675) 0.725 (0.736) 0.739 (0.772)
Stay at Home 0.623 (0.637) 0.672 (0.691) 0.669 (0.681) 0.699 (0.704) 0.675 (0.662) 0.705 (0.711) 0.725 (0.727)
Average Performance 0.579 (0.589) 0.728 (0.740) 0.674 (0.665) 0.735 (0.734) 0.704 (0.685) 0.729 (0.712) 0.756 (0.745)
V-A Baselines
We consider the following baselines. Table II shows the difference between baselines and the proposed approach. Notably, we have implemented baselines for both cross-domain and cross-target tasks regardless of their specific task targets. This decision stems from the unique nature of the task, which presents limitations in terms of available baselines. We chose BERT as our base model instead of larger language models due to its demonstrated effectiveness in text classification, accessibility to task-specific pretrained models, and computational efficiency.
RLFND [41]: A domain-adaptive reinforcement learning framework for fake news detection. The proposed method leverages a deep reinforcement learning algorithm to learn the optimal policy for fake news detection while adapting to the domain shift between the source and target domains. MTSD [24]: A multi-task framework that performs stance detection on multiple targets. The framework leverages a shared encoder to capture the common features of the text and a task-specific decoder to predict the stance toward each target.
Llama-2-7B [42]: A pre-trained and fine-tuned large language model released by Meta AI, designed for a variety of NLP tasks. We use the smallest variant for our problem setting. Due to Llama-2’s incapability of performing stance detection on our dataset with an acceptable performance, NeMo framework was used to perform Supervised Fine-Tuning (SFT) on this model.
V-B Datasets
We use two representative datasets related to COVID-19, each from different domains and targets (details provided in Table I). The large-scale Twitter dataset CoVaxNet [43] was employed for training across both domains and targets. CoVaxNet was compiled throughout the pandemic period (Jan 1, 2020 - Dec 31, 2021) and annotated with stances towards COVID-19 vaccination. To evaluate the performance of STANCE-C3  in a single-target context, we categorized the domains as distinct time intervals distinguished by significant events, following common practice in social media analysis [44]. We divided CoVaxNet into two periods, with the full approval of the Pfizer-BioNTech vaccine by the US Food and Drug Administration (FDA) on Aug 23, 2021, serving as the dividing event. Thus, the source and target domains consist of 10,000 tweets each, selected randomly from January 1, 2020 to Aug 22, 2021 (“CoVaxNet_pre”; source domain) and from Aug 23, 2021 to Dec 31, 2021 (“CoVaxNet_post”; target domain).
For the cross-target scenario, CoVaxNet_pre and CoVaxNet_post were separately used as training sets for STANCE-C3. Model performance was evaluated on the targets (“Anthony S. Fauci, M.D.”, “Wearing a Face Mask”, “Keeping Schools Closed”, “Stay at Home Orders”) from the COVID-19-Stance [45] dataset [^2] Dataset collected between Feb 27, 2020 and Aug 20, 2020. It includes different targets and keywords of interest (e.g., #lockdown), thereby justifying an adequate distributional shift between the source and target. Domain counterfactuals for this dataset were generated in the same manner as the cross-domain setting to augment the training set. Importantly, we labeled the generated examples the same as the original example as we observed a positive correlation between the targets in the source and target domain. Future works should explore scenarios where the stances towards targets in source and target domains are conflicting.
We excluded tweets that were too short or contained excessive hashtags, emojis, and URLs from both datasets to preserve their quality. Additionally, we balanced the training and test set across the different labels. It should be noted that congruent labels from the two datasets were treated as equivalent; that is, pro is equivalent to in-favor and anti is the same as against. In the context of stance detection, the terms in these pairs essentially convey the same stance. Both CoVaxNet and COVID-19-Stance datasets are publicly available for research purposes [43].
V-C Evaluation and Results
Table III and Table IV summarize the comparative performance of the baseline and proposed models concerning the accuracy and AUC (in parentheses).
Q1: In this experiment, the proposed approach and the domain-adaptive baselines used a fixed portion (30%) of the target domain data during counterfactual data generation and maintained the same ratio in the training set. The models were trained on a single domain and tested on both source and target domains for stance detection for a single target of interest from the source domain. Our results provide some evidence that even when the target of interest is contained in the source domain, performance in the target domain generally degrades for all models. The Performance Degradation in the table highlights how a model’s performance changes when shifting from the source domain to the target domain. We observe that all models exhibit performance degradation, confirming the difference between domains.
Q2: The results of our experiments on the cross-domain cross-target setting, as presented in Table IV, demonstrate the performance of our approach and the baselines. In this setting, we augment the training set of baselines with a small portion (30%) of data from the target domain. The results of our experiments suggest that the proposed approach is more effective than the domain-adaptive baselines. Specifically, our approach surpasses all baselines regarding average accuracy and AUC score, underscoring its suitability for the cross-domain cross-target context. Furthermore, these results underscore the robustness of our approach to variations in the target domain, a strength attributable to our counterfactual data generation technique which is tailored to produce data that better represents the target domain.
Parameter Analysis To address Q3, we analyze the impact of hyperparameters on the performance of STANCE-C3. To demonstrate the effect of contrastive loss and the balance between contrastive and cross-entropy loss, we vary the $\lambda$ value in the range of $\{0.0,0.25,0.5,0.75,1.0\}$ during training. As illustrated in Figure 3-a, the AUC score varies significantly for different $\lambda$ values. The results suggest that using $\lambda=0.5$ yields the best performance.
We further investigate the impact of the target domain data portion $\gamma$ utilized for counterfactual data generation by varying its value within the range of $\{5\%,15\%,30\%,45\%\}$. The results, as illustrated in Figure 3-b, demonstrate that incorporating a larger portion of target domain data leads to enhanced performance. However, in practical scenarios, access to substantial data in the target domain is often constrained. Considering there is no statistically significant performance difference observed between $\gamma=30\%$ and $\gamma=45\%$, we conclude that employing 30% of the target domain data enables an acceptable level of performance.